# Smooth Tests for Normality in ANOVA

### v3

@@ 3.3 Same mean and different variances
where $\{\sigma_{j}\}_{j=1}^{J}$ represents the potentially distinct standard deviations of each group. For estimation, we adopt $\hat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}(Y_{ij}-\hat{\mu})^{2}$ for $\mu$ and $\sigma_{j}^{2}$ ($j=1,\ldots,J$), respectively. Note that the estimator of the population mean here is different from the sample mean in Section 3.1 due to the heterogeneity of group variances. Define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu})/\hat{\sigma}_{j}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. Let $p_{j}=\lim N_{j}/N$ and $q_{j}=\lim JN_{j}/N$ be quantities characterizing the relative proportions of the group sample sizes. Here we further impose the following conditions for model (3.10): (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$. These conditions imply that the group-specific sample sizes and standard deviations are of the same order of magnitude.
Following the approaches in Sections 3.1 and 3.2, we first examine the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, which is summarized in the theorem below.
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.11)
@@
Note that the asymptotic decomposition of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ in Theorem 7 is no longer the same as that in Theorems 1 and 4. The difference lies in the parameter estimation effect of $\hat{\mu}$, which is reflected by the term $c_{1k}N^{-1}(\sum_{\ell=1}^{J}p_{\ell}\sigma^{-1}_{\ell})\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\sigma_{j}e_{ij}/q_{j}$ on the right-hand side of (3.11). Unlike model (3.1), which uses the sample mean to estimate $\mu$, model (3.10) employs a different estimator, $\hat{\mu}$, resulting in this distinct estimation effect. In particular, when all groups share the same variance, i.e., $\sigma_{1}=\ldots=\sigma_{J}$, the term reduces to $c_{1k}N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}=-c_{1k}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}$; further if the group sample sizes are equal, i.e., $N_{1}=\ldots=N_{J}$, then the estimation effect simplifies to that in Theorem 1, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$. In contrast, the estimation effect of the variances always aggregates into the term $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$ regardless of the equality of group variances or group sample sizes.
The asymptotic normality of the vector $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$ follows directly from (3.11). Under the conditions of Theorem 7, an application of the CLT yields that
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$
$\to_{d}\sum_{j=1}^{J}\sqrt{p_{j}}\bm{Z}_{j},$
@@
$\omega_{kl}^{(j)}$ $=\mathrm{E}\left\{\pi_{k}(Z)-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{c_{2k}}{2}\left(e^{2}-1\right)\right\}$
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\to_{d}\mathcal{N}_{K}\left(\bm{0},\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right).$
By replacing $\sigma_{j}$, $p_{j}$ and $q_{j}$ with their sample analogues $\hat{\sigma}_{j}$, $\hat{p}_{j}=N_{j}/N$ and $\hat{q}_{j}=JN_{j}/N$ respectively, we obtain a consistent estimator of $\bm{\Omega}_{K}^{(j)}$, denoted by $\widehat{\bm{\Omega}}_{K}^{(j)}$. Accordingly, the test statistic is defined as
$\hat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}\hat{p}_{j}\widehat{\bm{\Omega}}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.12)
@@
$\tilde{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.13)
@@
Corollary 2 shows that under the null hypothesis, both the feasible test statistic $\hat{\Psi}_{K}^{2}$ (defined in (3.12)) and its infeasible counterpart $\tilde{\Psi}_{K}^{2}$ (in (3.13)) are asymptotically $\chi_{K}^{2}$ distributed when multiplied by the total sample size $N$. Although these test statistics take a slightly different form from those in the previous Sections 3.1 and 3.2, the limiting Chi-square distribution is retained, and the convergence rate remains $N$, which further supports the validity of our unified theoretical framework. The following two theorems characterize the asymptotic behavior of $\tilde{\Psi}_{K}^{2}$ under both fixed and local alternatives, in a manner analogous to that of $\hat{\Psi}_{K}^{2}$ discussed in Sections 3.1 and 3.2.
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}\right$.
$-\left.\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (3.14)
@@
$\tilde{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K},$
@@
$\sqrt{N}\left(\tilde{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right)\to_{d}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Upsilon}_{K}\bm{a}_{K}\right),$ (3.15)
@@
$\lambda_{kl}^{(j)}=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\frac{(d_{1k}d_{3l}+d_{1l}d_{3k})\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{d_{1k}d_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}$
$+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{(d_{1k}d_{2l}+d_{1l}d_{2k})\sigma_{j}}{2q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\mathrm{E}\left[e^{3}\right]$
$+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}\right$.
@@
$N\tilde{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{\Delta}_{K}\right).$
@@
To conclude this section, we summarize the regularity conditions for the three models in Table 1. On one hand, more complex model structures naturally impose additional restrictions on parameters and group sample sizes, as the heterogeneity across groups introduces extra challenges for theoretical analysis. On the other hand, under these mild conditions, an asymptotic Chi-square test with convergence rate $N$ can be established for each model. This highlights that our proposed methodology provides a unified theoretical framework while remaining flexible enough to accommodate the distinctive features of different ANOVA settings.
Table 1: The regularity conditions for the three models
@@
Same mean & same variance (3.1) $N\to\infty$
Different means & same variance (3.9)
$\min\{N_{1},\ldots,N_{J}\}\to\infty$,
$J=o(N^{1/2})$, $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$
Same mean & different variances (3.10)
$\min\{N_{1},\ldots,N_{J}\}\to\infty$,
$J=o(N^{1/2})$ , $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$,
$\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$,
$\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$

### v4

@@ 3.3 Common group means with heterogeneous variances
where $\{\sigma_{j}\}_{j=1}^{J}$ represents the potentially distinct standard deviations of each group. For estimation, we adopt $\widehat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\widehat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}(Y_{ij}-\widehat{\mu})^{2}$ for $\mu$ and $\sigma_{j}^{2}$ ($j=1,\ldots,J$), respectively. Note that the estimator of the population mean here is different from the sample mean in Section 3.1 due to the heterogeneity of group variances. Define $\widehat{\varepsilon}_{ij}=Y_{ij}-\widehat{\mu}$, $\widehat{e}_{ij}=\widehat{\varepsilon}_{ij}/\widehat{\sigma}_{j}=(Y_{ij}-\widehat{\mu})/\widehat{\sigma}_{j}$ and $\widehat{Z}_{ij}=\Phi(\widehat{e}_{ij})$. Let $p_{j}=\lim N_{j}/N$ and $q_{j}=\lim JN_{j}/N$ be quantities characterizing the relative proportions of the group sample sizes. Here we further impose the following conditions for model (3.10): (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$. These conditions imply that the group-specific sample sizes and standard deviations are of the same order of magnitude.
Following the approach in Sections 3.1 and 3.2, we first examine the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$, which is summarized in the theorem below.
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.11)
@@
Note that the asymptotic decomposition of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ in Theorem 7 is no longer the same as that in Theorems 1 and 4. The difference lies in the parameter estimation effect of $\widehat{\mu}$, which is reflected by the term $c_{1k}N^{-1}(\sum_{\ell=1}^{J}p_{\ell}\sigma^{-1}_{\ell})\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\sigma_{j}e_{ij}/q_{j}$ on the right-hand side of (3.11). Unlike model (3.1), which uses the sample mean to estimate $\mu$, model (3.10) employs a different estimator, $\widehat{\mu}$, resulting in this distinct estimation effect. In particular, when all groups share the common variance, i.e., $\sigma_{1}=\ldots=\sigma_{J}$, the term reduces to $c_{1k}N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}=-c_{1k}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}$; further if the group sample sizes are equal, i.e., $N_{1}=\ldots=N_{J}$, then the estimation effect simplifies to that in Theorem 1, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$. In contrast, the estimation effect of the variances always aggregates into the term $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$ regardless of the equality of group variances or group sample sizes.
The asymptotic normality of the vector $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\widehat{Z}_{ij})$ follows directly from (3.11). Under the conditions of Theorem 7, an application of the CLT yields that
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\widehat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\widehat{Z}_{ij}\right)\right]^{\top}$
$\overset{d}{\to}\sum_{j=1}^{J}\sqrt{p_{j}}\bm{Z}_{j},$
@@
$\omega_{kl}^{(j)}$ $=\mathbb{E}\left\{\pi_{k}(Z)-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{c_{2k}}{2}\left(e^{2}-1\right)\right\}$
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\overset{d}{\to}\mathcal{N}_{K}\left(\bm{0},\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right).$
By replacing $\sigma_{j}$, $p_{j}$ and $q_{j}$ with their sample analogues $\widehat{\sigma}_{j}$, $\widehat{p}_{j}=N_{j}/N$ and $\widehat{q}_{j}=JN_{j}/N$ respectively, we obtain a consistent estimator of $\bm{\Omega}_{K}^{(j)}$, denoted by $\widehat{\bm{\Omega}}_{K}^{(j)}$. Accordingly, the test statistic is defined as
$\widehat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}\widehat{p}_{j}\widehat{\bm{\Omega}}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right).$ (3.12)
@@
$\widetilde{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)\right).$ (3.13)
@@
Corollary 2 shows that under the null hypothesis, both the feasible test statistic $\widehat{\Psi}_{K}^{2}$ (defined in (3.12)) and its infeasible counterpart $\widetilde{\Psi}_{K}^{2}$ (in (3.13)) are asymptotically $\chi_{K}^{2}$ distributed when multiplied by the total sample size $N$. Although these test statistics take a slightly different form from those in the previous Sections 3.1 and 3.2, the limiting Chi-square distribution is retained, and the convergence rate remains $N$, which further supports the validity of our unified theoretical framework. The following two theorems characterize the asymptotic behavior of $\widetilde{\Psi}_{K}^{2}$ under both fixed and local alternatives, in a manner analogous to that of $\widehat{\Psi}_{K}^{2}$ discussed in Sections 3.1 and 3.2.
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}\right$.
$-\left.\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+\mathbb{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (3.14)
@@
$\widetilde{\Psi}_{K}^{2}\overset{p}{\to}\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K},$
@@
$\sqrt{N}\left(\widetilde{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right)\overset{d}{\to}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Upsilon}_{K}\bm{a}_{K}\right),$ (3.15)
@@
$\lambda_{kl}^{(j)}=$ $\mathbb{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\frac{(d_{1k}d_{3l}+d_{1l}d_{3k})\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{d_{1k}d_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}$
$+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{(d_{1k}d_{2l}+d_{1l}d_{2k})\sigma_{j}}{2q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\mathbb{E}\left[e^{3}\right]$
$+\frac{d_{2k}d_{2l}}{4}\left[\mathbb{E}\left[e^{4}\right]-1\right].$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}\right$.
@@
$N\widetilde{\Psi}_{K}^{2}\overset{d}{\to}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{\Delta}_{K}\right).$
@@
To conclude this section, we summarize the regularity conditions on sample size structure and cross-group heterogeneity for the three models (3.1), (3.9), and (3.10) in Table 1. On the one hand, more complex model structures naturally impose additional restrictions on within-group variability and group sample sizes, as heterogeneity across groups introduces additional challenges for theoretical analysis. On the other hand, under these mild conditions, an asymptotic Chi-square test with convergence rate $N$ can be established for each model. This highlights that our proposed methodology provides a unified theoretical framework while remaining flexible enough to accommodate the distinctive features of various ANOVA settings.
Table 1: Regularity conditions on sample sizes and group heterogeneity for the three one-way models
@@
Model (3.1) $N\to\infty$
Model (3.9) $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$, $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$
Model (3.10) $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ , $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$, $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$, $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$